the log base is in parenteses right after the log
Log(2) (x-8) + Log(2) (x+4) = 5

Answer by KnightOwlTutor(293)

Change form to exponent
(x-8)(x+4=2^5=32
Use FOIL to ge x^2-4x=0
factorx(x-4)=0
x=0 or 4
In logs you only use + numbers
Therefore 4 is the correct answer.